OpenStudy (anonymous):

for the function f(x)=(8-2x)^2, find f^-1,determine whether f^-1 is a function HELP PLEASE :))

OpenStudy (anonymous):

inverse is 

OpenStudy (anonymous):

To find the inverse of a function swap the positions of x and y. Pretend that F(x) is y, so y = (8 - 2x)^2 If we swap them it becomes. x = (8 - 2y)^2 Now we solve for y sqrt(x) = 8 - 2y sqrt(x) - 8 = -2y y = (8 - sqrt(x)) / 2

OpenStudy (anonymous):

is it a function?

OpenStudy (anonymous):

Yes, it is a function.

OpenStudy (anonymous):

As long as x is never negative, however

OpenStudy (anonymous):

thank you SO much, i really do appreciate the help (:

OpenStudy (anonymous):

No problem XD.